OpenStudy (anonymous):

Please Help!!!!!!!!!!!!!!!!!!!!!!!!!!!! Solve the equations by substitution. If the equations are inconsistent and the solution is an empty set, write "no solution". 2x + y = 7 y = x + 1 x= y=

OpenStudy (anonymous):

@OrangeMaster

OpenStudy (anonymous):

I was taught this a year ago, but things that will be useless in life later, kinda get lost and forgotten in my brain.

OpenStudy (anonymous):

substitution is when you substitute one equation into another to find your unknown 2x+y=7 --- equation (1) y= x+1 --- equation (2) sub (2) into (1) by 2x + (x+1) = 7 now solve for x

OpenStudy (anonymous):

Sorry that I can't help you.

OpenStudy (anonymous):

2x+x+1=7 find x and then y from any equation.

OpenStudy (anonymous):

is it not clear how to solve \[2x+x+1=7\]?